OpenStudy (anonymous):

(2x^2+6x+1)+(-7x^2 +2x - 3)

OpenStudy (anonymous):

\[\left(-7 x^2+2 x-3\right)+\left(2 x^2+6 x+1\right)=-5 x^2+8 x-2 \]